# Exactly solvable dissipative dynamics and one-form strong-to-weak spontaneous symmetry breaking in interacting two-dimensional spin systems

### v1

@@ Appendix A Nontrivial steady states: no-go result
$L\rho L^{\dagger}=\sum_{ij,\mu\nu}\ell_{i,\mu}\ell^{*}_{j,\nu}v_{ij}^{\mu\nu}\mathrm{i}c_{i}\tilde{c}_{j},$ (67)
where, analogously to before, $v_{ij}^{\mu\nu}=\mathrm{i}b^{\mu}_{i}\tilde{b}_{j}^{\nu}$. Two operators $v_{ij}^{\mu\nu}$ anticommute whenever they share an odd number of indices (they also anticommute with certain $u_{ij}$ and $\tilde{u}_{ij}$ that arise in the Hamiltonian contribution). Since the full set of edge operators $\{u_{ij},\tilde{u}_{ij},v_{ij}^{\mu\nu}\}$ does not commute with itself, the operators cannot be replaced by their eigenvalues ±$\pm 1$ and the fermionized Lindbladian does not split into individual free-fermion sectors. This only occurs when each $L$ is a single gamma matrix, whence a trivial steady state follows.
[…]
Figure 8: Additional numerical results for the spectral gap as a function of dissipation strength in sectors with fermion-number conservation.‘ We computed the gap for $n=1$ on the honeycomb lattice with $N=512$ (top row) and triangular lattice with $N=400$ (bottom row) for periodic boundary conditions on the torus, using the prescription of Eq. (64). We set $\gamma_{j}=\gamma X_{j}$ with $X_{j}$ independently sampled from a uniform distribution in $[0,2]$ for each site $j$ and disordered couplings $J_{ij}$ independently sampled from $[-2,2]$. The left panels correspond to a single interlayer gauge flip ($|\mathfrak{V}|=1$) and the right panels to three flips ($|\mathfrak{V}|=3$). In each panel, we show the results for the different background flux configurations described in the caption of Fig. 6. The shaded gray area gives the analytical bounds (65), which, as expected, are less tight than in the uniform case.
@@ Appendix B Counting of independent fluxes
First, we consider the fluxes through the elementary plaquettes $p$ of the graph ($W_{p}$ and $\widetilde{W}_{p}$), of which there are $F$ in each copy. Additionally, for each copy, we can have $b_{1}$ noncontractible loops $\ell$ and the associated fluxes $W_{\ell}$ and $\widetilde{W}_{\ell}$. There is also a weak flux $\mathcal{W}_{q}$ through each interlayer plaquette $q$, of which is there one for each of the $E$ edges. In total, we thus have $2(F+b_{1})+E$ elementary fluxes.
The flux through any closed surface is $+1$ because each gauge field, which squares to $+1$, appears twice in the definition of the flux through the surface. Then, if $S$ is closed (i.e., $b_{2}=1$), the total flux through the surface (i.e., $\prod_{p}W_{p}=+1$) is fixed. Moreover, the flux through each of the $F$ interlayer cells (i.e., $W_{p}\widetilde{W}_{p}\prod_{q\in\partial p}\mathcal{W}_{q}=+1$) fixes all the fluxes $\widetilde{W}_{p}$ since there is one $\widetilde{W}_{p}$ per interlayer cell. (The flux $\prod_{p}\widetilde{W}_{p}$ is already fixed by the previous conditions and, therefore, does not amount to an additional constraint.) Finally, $W_{\ell}\widetilde{W}_{\ell}=\prod_{q}\mathcal{W}_{q}$, where the product is taken over interlayer plaquettes $q$ that connect the noncontractible loops $\ell$ in the two copies; this fixes all fluxes $\widetilde{W}_{\ell}$ and gives $b_{1}$ further constraints.
In total, we thus have $F+b_{1}+b_{2}$ constraints, and the number of independent fluxes (gauge-invariant degrees of freedom) is $F+E+b_{1}-b_{2}$: $F-b_{2}$ plaquette fluxes $W_{p}$, $b_{1}$ noncontractible-loop fluxes $W_{\ell}$, and $E$ interlayer weak fluxes $\mathcal{W}_{q}$, in agreement with Eq. (43). Using the Euler characteristic, the number of independent fluxes can be written as $12E-N+$ 1.
@@ Appendix D Bendixson inequalities
Let $M$ be an arbitrary real matrix, $z$ any one of its eigenvalues, and $x_{m}$ and $x_{M}$ respectively the smallest and largest eigenvalues of $(M+M^{T})/2$, where ${M}^{T}$ denotes the transpose of $M$. Then,

### v2

@@ Acknowledgements.
$L\rho L^{\dagger}\to\sum_{ij,\mu\nu}\ell_{i,\mu}\ell^{*}_{j,\nu}v_{ij}^{\mu\nu}\mathrm{i}c_{i}\tilde{c}_{j}|\rho\rangle,$ (68)
where, analogously to before, $v_{ij}^{\mu\nu}=\mathrm{i}b^{\mu}_{i}\tilde{b}_{j}^{\nu}$. Two operators $v_{ij}^{\mu\nu}$ anticommute whenever they share an odd number of indices (they also anticommute with certain $u_{ij}$ and $\tilde{u}_{ij}$ that arise in the Hamiltonian contribution). Since the full set of edge operators $\{u_{ij},\tilde{u}_{ij},v_{ij}^{\mu\nu}\}$ does not commute with itself, the operators cannot be replaced by their eigenvalues $\pm 1$, and the fermionized Lindbladian does not split into individual free-fermion sectors. This only occurs when each $L$ is a single gamma matrix, whence a trivial steady state follows.
Figure 7: Additional numerical results for the spectral gap as a function of dissipation strength in sectors with fermion-parity conservation. The gap was computed for the triangular lattice with N=144N=144576576 and periodic boundary conditions on the torus, using the prescription of Sec. V.3. Different panels correspond to the different choices of intralayer gauge flips (i.e., of $\mathfrak{U}$) described in the caption of Fig. 5.
Figure 8: Additional numerical results for the spectral gap as a function of dissipation strength in sectors with fermion-number conservation.‘ We computed the gap for $n=1$ on the honeycomb lattice with $N=512$ (top row) and triangular lattice with $N=400$ (bottom row) for periodic boundary conditions on the torus, using the prescription of Eq. (65). We set $\gamma_{j}=\gamma X_{j}$ with $X_{j}$ independently sampled from a uniform distribution in $[0,2]$ for each site $j$ and disordered couplings $J_{ij}$ independently sampled from [2,2][-2,2]. The left panels correspond to a single interlayer gauge flip ($|\mathfrak{V}|=1$) and the right panels to three flips ($|\mathfrak{V}|=3$). In each panel, we show the results for the different background flux configurations described in the caption of Fig. 6. The shaded gray area gives the analytical bounds (66), which, as expected, are less tight than in the uniform case.
@@
First, we consider the fluxes through the elementary plaquettes $p$ of the graph ($W_{p}$ and $\widetilde{W}_{p}$), of which there are $F$ in each copy. Additionally, for each copy, we can have $b_{1}$ noncontractible loops $\ell$ and the associated fluxes $W_{\ell}$ and $\widetilde{W}_{\ell}$. There is also a weak flux $\mathcal{W}_{q}$ through each interlayer plaquette $q$, of which there is one for each of the $E$ edges. In total, we thus have $2(F+b_{1})+E$ elementary fluxes.
The flux through any closed surface is +1+1 because each gauge field, which squares to +1+1, appears twice in the definition of the flux through the surface. Then, if $S$ is closed (i.e., $b_{2}=1$), the total flux through the surface (i.e., $\prod_{p}W_{p}=+1$) is fixed. Moreover, the flux through each of the $F$ interlayer cells (i.e., $W_{p}\widetilde{W}_{p}\prod_{q\in\partial p}\mathcal{W}_{q}=+1$) fixes all the fluxes $\widetilde{W}_{p}$ since there is one $\widetilde{W}_{p}$ per interlayer cell. (The flux $\prod_{p}\widetilde{W}_{p}$ is already fixed by the previous conditions and, therefore, does not amount to an additional constraint.) Finally, $W_{\ell}\widetilde{W}_{\ell}=\prod_{q}\mathcal{W}_{q}$, where the product is taken over interlayer plaquettes $q$ that connect the noncontractible loops $\ell$ in the two copies; this fixes all fluxes $\widetilde{W}_{\ell}$ and gives $b_{1}$ further constraints.
In total, we thus have $F+b_{1}+b_{2}$ constraints, and the number of independent fluxes (gauge-invariant degrees of freedom) is $F+E+b_{1}-b_{2}$: $F-b_{2}$ plaquette fluxes $W_{p}$, $b_{1}$ noncontractible-loop fluxes $W_{\ell}$, and $E$ interlayer weak fluxes $\mathcal{W}_{q}$, in agreement with Eq. (43). Using the Euler characteristic, the number of independent fluxes can be written as 2EN+12E-N+1.
@@
Let $M$ be an arbitrary real matrix, $z$ any one of its eigenvalues, and $x_{m}$ and $x_{M}$, respectively, the smallest and largest eigenvalues of $(M+M^{T})/2$, where ${M}^{T}$ denotes the transpose of $M$. Then,